# Maritime Cybersecurity: A Comprehensive Review

### v2

@@ Abstract.
Maritime cyberattacks, maritime cybersecurity, maritime incidents, operational technology system, vulnerabilities, threats, countermeasures
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 11journalnumber: 1article: 111publicationmonth: 1ccs: General and reference Surveys and overviewsccs: Security and privacy Systems security
1. Introduction
@@
The global maritime digital technology industry is estimated to be worth \$345 billion by 2030, up from a previous forecast of \$279 billion (International, 2021). Over the past few years, the maritime industry has progressed rapidly and its evolution has extended into several domains. From increased use of digital systems (Gavalas et al., 2022), enhanced satellite communication (Alqurashi et al., 2022), and Internet-of-Things (IoT)-enabled port infrastructure (Cil et al., 2022) to development tendencies of autonomous shipping (Karetnikov et al., 2019), a series of studies have been conducted with a focus on the advancements of maritime technology. Despite this progress, maritime cybersecurity has been inadequately addressed. Specifically, the maritime industry is at risk of becoming a target for a series of cyber threats and attack vectors that arise from the interconnection of diverse technical tools. This includes, among others, IoT networks and telecommunications. The maritime cyber environment, as described by Sotiria (Lagouvardou, 2018), encompasses the networked systems of both Information Technology (IT) and cyber-physical systems (i.e., Operational Technology (OT)). It includes communication networks that allow data to flow from IT systems of a ship to the OT systems. Such is accomplished via programmable logic controllers (PLCs) and various sensors and ultimately controls advanced navigation tools like Global Positioning Systems (GPS).
The integration of IT and OT in the maritime sector introduces heightened cybersecurity risks, as traditionally isolated OT systems become more susceptible to cyber attacks through their connections to IT networks (depicted in Figure 1). As the range of potential threats broaden, more vulnerabilities and entry points have been discovered in maritime systems.
@@
Figure 1. Illustration of a typical maritime shipboard IT/OT system that combines hardware and software to monitor and control physical devices in real time. The increasing interconnectivity of IT and OT systems have expanded the potential attack surface.
@@
Vessels have become vulnerable to cyberattacks originating from both internally and externally. With the rising levels of connectivity, the scope of maritime cybersecurity extends beyond the vessels and is characterized by complex interactions among stakeholders and components, such as shippers and ports. Unlike the automotive industry, where the impact of cyberattacks tend to be more localized, affecting single vehicles or restricted regions, the potential impact of a cyberattack on a large vessel could destabilize global supply chains and inflict billions in economic losses (Roberts et al., 2019). Moreover, the risk extends beyond individual ships - a cyberattack targeting shared infrastructure like a VSAT service supplier could trigger a fleetwide or cross-fleet event (FLINT, 2022), magnifying the damage across multiple vessels and shipping companies simultaneously.
In the last few years, several surveys on maritime cybersecurity have been published (Ashraf et al., 2022; Ben Farah et al., 2022, 2022; Park et al., 2019; Yu et al., 2023). Park et al. (Park et al., 2019) adopted a high-level perspective of cybersecurity threats and risk control alternatives in the marine industry as a whole. Ismail et al. (Ismail et al., 2021) focused specifically on maritime cybersecurity in the Indian Ocean region with limited depth on maritime component-specific topics. Farah et al. (Ben Farah et al., 2022) provided a mapping and classification of on-vessel core equipment and in-port infrastructure and included a detailed breakdown of electro-mechanical systems, electronic systems and communication systems. Nevertheless, the authors (Ben Farah et al., 2022) did not offer comprehensive technical information on particular cybersecurity measures. In contrast to more general maritime cybersecurity, Ashraf et al. (Ashraf et al., 2022) offered a comprehensive classification of cyberattacks in the marine domain, with a primary focus on risks related to the Internet of Things. However, this work also did not elaborate on methods of mitigation. Yu et al. (Yu et al., 2023) used bibliometric methods to provide a review of academic literature on maritime cybersecurity with limited discussion of technical details in cybersecurity solutions. Hence, there is still room to classify novel threats and countermeasures in literature that are more sophisticated and highly targeted, and our survey aims to fill this gap.
Although security attacks against maritime IoT devices have received considerable attention during the last decade, the significance of IoT enabled attacks is not always fully assessed. We focus on verified attacks, i.e, either real-world incidents or attacks that have been implemented and recently published by researchers including the attacks that are theoretically possible. Based on the analysis of these attacks, we summarize security solutions that can effectively mitigate such threats. Unlike existing works that involve limited discussions on countermeasures, our study will provide a thorough review of the latest defensive strategies and technologies. Concretely, we cover not only component-specific countermeasures but also integrated solutions that can enhance the overall cybersecurity posture of maritime organizations. Table 1 provides a comparison between our work and prior works targeting maritime cybersecurity.
@@
Table 1. Positioning of our study with respect to prior surveys on maritime cybersecurity
@@
  1. (1)
    We provide a broader overview of the industry’s cybersecurity landscape by examining six key maritime interconnected aspects in Section 3.
  2. (2)
    Section 4 reviews the threats of vessel components, which include RAdio Detection And Ranging (RADAR), Voyage Data Recorders (VDR), Global Navigation Satellite Systems (GNSS) / GPS, Electronic Chart Display and Information Systems (ECDIS), Automatic Identification Systems (AIS), Global Maritime Distress and Safety System (GMDSS) and Very Small Aperture Terminal (VSAT).
  3. (3)
    We illustrate and discuss a taxonomy that categorizes maritime incidents over a 10-year period, from 2014 to 2023, targeting various attack surfaces and systems of vessel in Section 4.
  4. (4)
    We provide a review of diverse mitigation techniques aimed at addressing security concerns in Section 5.
  5. (5)
    After we discuss challenges and potential future directions for maritime cybersecurity in Sections 5-6, we conclude our article in Section 7.
Figure 2. Structure of this study
2. Literature Search Methodology
In this section, we will discuss our research protocol in selecting and investigating the prior literature for our survey article.
2.1. Keywords and Databases
To narrow the scope of this survey and ensure its relevance to the maritime industry and academic research, a comprehensive set of keywords was meticulously chosen. The primary focus is on ”shipboard operational technology (OT) systems”, ”maritime cybersecurity”, and ”maritime industry”. These keywords are often used in conjunction with terms such as ”cyber threats”, ”vulnerabilities”, ”risks”, ”mitigation strategies”, ”countermeasures”, and ”cyber risk management” to identify the most pertinent literature addressing security challenges, potential risks, and defensive measures within the maritime cybersecurity domain. The following academic research databases were considered: Elsevier ScienceDirect, Association for Computing Machinery (ACM) digital library, IEEE Xplore digital library, and SpringerLink.
2.2. Literature Selection
While the selected scholarly databases are known for publishing high-quality research, we found that there was a limited number of studies specifically addressing the security of shipboard OT systems and the maritime industry as a whole. A reason could be that maritime cybersecurity is a niche area that intersects with various domains, and researchers may prefer to publish their work in domain-specific journals and conferences catering to the maritime industry.
To overcome this limitation, the scope was expanded to include non-traditional sources such as technical reports, white papers, and dissertations from reputable maritime associations, societies, and related organizations. These include guidelines and recommendations from the International Maritime Organization (IMO), the American Bureau of Shipping (ABS), the Baltic and International Maritime Council (BIMCO) and others, which complement the findings from academic publications. Additionally, maritime-specific journals such as The Journal of Navigation, published by Cambridge University Press, and the Journal of Marine Science and Engineering, an open-access journal by MDPI, were consulted to capture the latest developments and research in the field.
3. Challenges in Maritime Cybersecurity
@@
Figure 3. Six challenges in maritime cybersecurity (a) and the connection between them (b)
3.1. Vessel Security
@@
Vessel security forms the cornerstone of maritime cybersecurity, interacting with and influencing other critical aspects of the maritime network. Crew members are on the front lines of implementing security measures and are often the first to detect and respond to security threats. Well-trained crew members who are aware of security protocols and potential threats can significantly enhance a vessel’s security posture (Berg et al., 2013). However, human factors can introduce vulnerabilities (e.g. neglect security procedures). Vessel security is also governed by a complex web of international laws, conventions, and industry standards which set out security-related requirements for ships, ports, and government agencies. Although most marine incidents resulted from failing to follow international regulations related to maritime safety (Abuelenin, 2017), compliance with these regulatory frameworks should be viewed as a minimum standard rather than the ultimate goal of security efforts. When a vessel enters a port, it becomes subject to the port’s security procedures and infrastructure. The port security management measures can refer to the subsequent port security section (Section 3.2). The increasing digitization of port operations, including the use of automated systems for cargo handling and vessel traffic management, creates new cybersecurity considerations for vessels (Inkinen et al., 2019). Risk assessment methodology aims help systematically identify potential security threats to a vessel and includes the protection of human life and health (Papageorgiou et al., 2024). Based on the identified risks, these methods guide the development of appropriate security measures and contingency plans. In global supply chains, transit times are heavily influenced by vessel speed and routes. Vessel operations are generally under significant time pressure, aligning with the ”just-in-time” philosophy underlying most modern supply chains (Lind et al., 2021). This pressure can cause hasty security decisions, which could jeopardize vessel safety by emphasizing temporary solutions over long-term security solutions. Since many vessel components are integrated systems made by multiple companies that operate in different industries, producers must have a deep awareness of the hazards associated with the supply chain (Cho et al., 2022).
@@
This study examines the security of a particular part of the Integrated Bridge System (IBS), which could be an entry point of attack for ships. The security of the IBS is crucial as it represents a central point of control for the vessel. A vast network of systems and components is used by ships to function well in the challenging marine environment. These include navigation systems, communication equipment, and various sensors and actuators distributed throughout the ship. Better efficiency and operational capabilities are made possible by this integration, but there is a risk as well: if one component is hacked, it may give the attacker access to all other systems (Goodrum et al., 2018). Notably, Section 4 of this document delves into specific attacks targeting these systems and components with the potential consequences of such attacks.
3.2. Port Security
@@
With forecasts showing a strong compound annual growth rate of 7.3% from 2023 to 2031, the worldwide port security market is expected to increase significantly and could reach an estimated value of \$173.59 billion by 2031 (Intelligence, 2024). This trajectory reflects the continued development of security solutions to counter new threats in marine commerce, as well as the growing acknowledgment of port security as an essential part of the infrastructure of global trade. When port security is applied comprehensively, it includes a wide range of policies, practices, and systems that are put in place to protect port operations, infrastructure, and surrounding areas against a variety of threats, such as theft, smuggling, terrorism, and other criminal activity (Sergi and Reid, 2021). In order to combat the growing threat of combined cyber-physical attacks for ports, the future of port security is probably going to center on merging cyber and physical security measures (Adams et al., 2020). Its multifaceted nature highlights its complexity, necessitating the combination of technological, procedural, and physical protections to build a strong defense against both traditional and emerging threats.
@@
Numerous studies have addressed port infrastructure risk in detail (Kapalidis et al., 2022)(Progoulakis et al., 2023)(Progoulakis et al., 2021). Port security management plays a vital role in reducing a wide range of risks and threats that are present in maritime environments. This strategy includes a number of essential elements to protect maritime operations and infrastructure, including emergency response plans, surveillance and monitoring, perimeter security, access control systems, and cargo screening (Christopher, 2009). While access control systems use sophisticated identification credentials, biometric technologies (such as fingerprint scanning and facial recognition), and screening procedures to regulate entry to sensitive areas, perimeter security uses cutting-edge fencing systems, barriers, and intrusion detection technologies to create a secure boundary around port facilities (Kusi, 2015). Furthermore, state-of-the-art monitoring systems for ship movements in port waters, along with non-intrusive cargo inspection methods like X-ray scanners and RFID tracking for cargo integrity, support prompt response times to security incidents (Christopher, 2009). The integrity, safety, and operational continuity of port infrastructures—which are crucial hubs in both national security frameworks and international trade networks—must be preserved by the integration of these security measures.
3.3. Supply Chain Security
@@
Supply chain security refers to the laws, practices, and technological advancements created to protect physical items, information flows, and supply chain assets against threats such as theft, terrorism, and natural catastrophes(Yang, 2011). The process of moving commodities by sea, involving shipping, port operations, and frequently interior transportation, is referred to as a maritime supply chain. This intricate system connects numerous ports spread over several nations and continents and encompasses a wide spectrum of stakeholders, including freight forwarders, shippers, carriers, and customs officials. The maritime supply chain is faced with a number of difficulties that could seriously jeopardize operational effectiveness and security. Among the many potential reasons of port interruptions, there include worker strikes, natural catastrophes, equipment malfunctions, and cyberattacks (Wendler-Bosco and Nicholson, 2020). Port disruptions are considered a critical vulnerability. Information security threats, such as cyberattacks on logistics management systems and data breaches revealing private shipping information, pose a significant concern as well (Liu et al., 2018). Physical security threats to marine supply chains include piracy (in high-risk maritime zones), cargo theft, and the smuggling of illegal commodities (Yang, 2011). These are problems that have not yet been properly resolved. As a result of continued climate change and rising sea levels, port infrastructure and shipping routes are becoming more and more susceptible to catastrophic weather events (Becker et al., 2013). Furthermore, operational risks such as equipment failures, delays in cargo handling, and human mistake in logistics planning can have a major impact on the efficiency of supply chains. Considering that over 80% of worldwide trade is carried out by ships (UNCTAD, 2022), any interruption to the movement of products around the world would be catastrophic.
@@
The suppliers’ perspective is a crucial aspect of the marine supply chain, particularly with regard to the tools and equipment they provide that are essential to numerous nautical systems. This aspect of supply chain security is vital and warrants detailed examination to fully understand the scope of maritime supply chain security. The maritime sector includes a range of specialized technology, including cargo handling devices, propulsion systems, communication tools, and numerous sensors. A breach in one or more product supply chains might lead to many vulnerabilities in important maritime systems. Malicious actors attempting to install hardware or software intended to interfere with operations and obtain unauthorized access may take advantage of these vulnerabilities (Liu et al., 2023). On the other hand, it may lead to the replacement of original parts with fake or inferior ones, which would raise the danger of equipment malfunctions and security issues (Yang, 2011). Moreover, the introduction of malicious code into software or firmware during the manufacturing process may result in the creation of a backdoor that allows for unauthorized control (Wendler-Bosco and Nicholson, 2020).
3.4. Risk Assessment Method

### v3

@@ Abstract
Maritime cyberattacks, maritime cybersecurity, maritime incidents, operational technology system, vulnerabilities, threats, countermeasures.
I Introduction
@@
The global maritime digital technology industry is estimated to be worth \$345 billion by 2030, up from a previous forecast of \$279 billion [1]. Over the past few years, the maritime industry has progressed rapidly and its evolution has extended into several domains. From increased use of digital systems [2], enhanced satellite communication [3], and Internet-of-Things (IoT)-enabled port infrastructure [4] to development tendencies of autonomous shipping [5], a series of studies have been conducted with a focus on the advancements of maritime technology. Despite this progress, maritime cybersecurity has been inadequately addressed. Specifically, the maritime industry is at risk of becoming a target for a series of cyber threats and attack vectors that arise from the interconnection of diverse technical tools. This includes, among others, IoT networks and telecommunications. The maritime cyber environment, as described by Sotiria [6], encompasses the networked systems of both Information Technology (IT) and cyber-physical systems (i.e., Operational Technology (OT)). It includes communication networks that allow data to flow from IT systems of a ship to the OT systems. Such is accomplished via programmable logic controllers (PLCs) and various sensors and ultimately controls advanced navigation tools like Global Positioning Systems (GPS). The integration of IT and OT in the maritime sector introduces heightened cybersecurity risks, as traditionally isolated OT systems become more susceptible to cyber attacks through their connections to IT networks (See Figure 1). As the range of potential threats broaden, more vulnerabilities and entry points have been discovered in maritime systems.
@@
Figure 1: Illustration of a typical maritime shipboard IT/OT system that combines hardware and software to monitor and control physical devices in real time. The increasing interconnectivity of IT and OT systems have expanded the potential attack surface.
@@
Vessels have become vulnerable to cyberattacks originating from both internally and externally. With the rising levels of connectivity, the scope of maritime cybersecurity extends beyond the vessels and is characterized by complex interactions among stakeholders and components, such as shippers and ports. Unlike the automotive industry, where the impact of cyberattacks tend to be more localized, affecting single vehicles or restricted regions, the potential impact of a cyberattack on a large vessel could destabilize global supply chains and inflict billions in economic losses [7]. Moreover, the risk extends beyond individual ships - a cyberattack targeting shared infrastructure like a VSAT service supplier could trigger a fleetwide or cross-fleet event [8], magnifying the damage across multiple vessels and shipping companies simultaneously.
@@
TABLE I: Positioning of our study with respect to prior surveys on maritime cybersecurity
@@
I-A Existing Work Comparisons
In the last few years, several surveys on maritime cybersecurity have been published [12, 11, 11, 9, 13]. Park et al. [9] adopted a high-level perspective of cybersecurity threats and risk control alternatives in the marine industry as a whole. Ismail et al. [10] focused specifically on maritime cybersecurity in the Indian Ocean region with limited depth on maritime component-specific topics. Farah et al. [11] provided a mapping and classification of on-vessel core equipment and in-port infrastructure and included a detailed breakdown of electro-mechanical systems, electronic systems and communication systems. Nevertheless, the authors [11] did not offer comprehensive technical information on particular cybersecurity measures. In contrast to more general maritime cybersecurity, Ashraf et al. [12] offered a comprehensive classification of cyberattacks in the marine domain, with a primary focus on risks related to the Internet of Things. However, this work also did not elaborate on methods of mitigation. Yu et al. [13] used bibliometric methods to provide a review of academic literature on maritime cybersecurity with limited discussion of technical details in cybersecurity solutions. Hence, there is still room to classify novel threats and countermeasures in literature that are more sophisticated and highly targeted, and our survey aims to fill this gap.
Although security attacks against maritime IoT devices have received considerable attention during the last decade, the significance of IoT enabled attacks is not always fully assessed. We focus on verified attacks, i.e, either real-world incidents or attacks that have been implemented and recently published by researchers including the attacks that are theoretically possible. Based on the analysis of these attacks, we summarize security solutions that can effectively mitigate such threats. Unlike existing works that involve limited discussions on countermeasures, our study will provide a thorough review of the latest defensive strategies and technologies. Concretely, we cover not only component-specific countermeasures but also integrated solutions that can enhance the overall cybersecurity posture of maritime organizations. Table I provides a comparison between our work and prior works targeting maritime cybersecurity.
I-B Our Contributions
@@
  1. 1.
    We provide a broader overview of the industry’s cybersecurity landscape by examining six key maritime interconnected aspects in Section II.
  2. 2.
    Section III-A reviews the threats of vessel components, which include RAdio Detection And Ranging (RADAR), Voyage Data Recorders (VDR), Global Navigation Satellite Systems (GNSS) / GPS, Electronic Chart Display and Information Systems (ECDIS), Automatic Identification Systems (AIS), Global Maritime Distress and Safety System (GMDSS) and Very Small Aperture Terminal (VSAT).
  3. 3.
    We illustrate and discuss a taxonomy that categorizes maritime incidents over a 10-year period, from 2014 to 2023, targeting various attack surfaces and systems of vessel in Section III-B.
  4. 4.
    We provide a review of diverse mitigation techniques aimed at addressing security concerns in Section IV-A and Section IV-B.
  5. 5.
    After we discuss challenges and potential future directions for maritime cybersecurity in Sections V-VI, we conclude our article in Section VII.
Figure 2: Structure of this study
II Challenges in Maritime Cybersecurity
@@
Figure 3: Six challenges in maritime cybersecurity (a) and the connection between them (b)
II-A Vessel Security
@@
Vessel security forms the cornerstone of maritime cybersecurity, interacting with and influencing other critical aspects of the maritime network. Crew members are on the front lines of implementing security measures and are often the first to detect and respond to security threats. Well-trained crew members who are aware of security protocols and potential threats can significantly enhance a vessel’s security posture [14]. However, human factors can introduce vulnerabilities (e.g. neglect security procedures). Vessel security is also governed by a complex web of international laws, conventions, and industry standards which set out security-related requirements for ships, ports, and government agencies. Although most marine incidents resulted from failing to follow international regulations related to maritime safety [15], compliance with these regulatory frameworks should be viewed as a minimum standard rather than the ultimate goal of security efforts. When a vessel enters a port, it becomes subject to the port’s security procedures and infrastructure. The port security management measures can refer to the subsequent port security section. The increasing digitization of port operations, including the use of automated systems for cargo handling and vessel traffic management, creates new cybersecurity considerations for vessels [16]. Risk assessment methodology aims help systematically identify potential security threats to a vessel and includes the protection of human life and health [17]. Based on the identified risks, these methods guide the development of appropriate security measures and contingency plans. In global supply chains, transit times are heavily influenced by vessel speed and routes. Vessel operations are generally under significant time pressure, aligning with the ”just-in-time” philosophy underlying most modern supply chains [18]. This pressure can cause hasty security decisions, which could jeopardize vessel safety by emphasizing temporary solutions over long-term security solutions. Since many vessel components are integrated systems made by multiple companies that operate in different industries, producers must have a deep awareness of the hazards associated with the supply chain [19].
@@
This study examines the security of a particular part of the Integrated Bridge System (IBS), which could be an entry point of attack for ships. The security of the IBS is crucial as it represents a central point of control for the vessel. A vast network of systems and components is used by ships to function well in the challenging marine environment. These include navigation systems, communication equipment, and various sensors and actuators distributed throughout the ship. Better efficiency and operational capabilities are made possible by this integration, but there is a risk as well: if one component is hacked, it may give the attacker access to all other systems [20]. Notably, Section 4 of this document delves into specific attacks targeting these systems and components with the potential consequences of such attacks.
II-B Port Security
@@
With forecasts showing a strong compound annual growth rate of 7.3% from 2023 to 2031, the worldwide port security market is expected to increase significantly and could reach an estimated value of \$173.59 billion by 2031 [21]. This trajectory reflects the continued development of security solutions to counter new threats in marine commerce, as well as the growing acknowledgment of port security as an essential part of the infrastructure of global trade. When port security is applied comprehensively, it includes a wide range of policies, practices, and systems that are put in place to protect port operations, infrastructure, and surrounding areas against a variety of threats, such as theft, smuggling, terrorism, and other criminal activity [22]. In order to combat the growing threat of combined cyber-physical attacks for ports, the future of port security is probably going to center on merging cyber and physical security measures [23]. Its multifaceted nature highlights its complexity, necessitating the combination of technological, procedural, and physical protections to build a strong defense against both traditional and emerging threats.
@@
Numerous studies have addressed port infrastructure risk in detail [24][25][26]. Port security management plays a vital role in reducing a wide range of risks and threats that are present in maritime environments. This strategy includes a number of essential elements to protect maritime operations and infrastructure, including emergency response plans, surveillance and monitoring, perimeter security, access control systems, and cargo screening [27]. While access control systems use sophisticated identification credentials, biometric technologies (such as fingerprint scanning and facial recognition), and screening procedures to regulate entry to sensitive areas, perimeter security uses cutting-edge fencing systems, barriers, and intrusion detection technologies to create a secure boundary around port facilities [28]. Furthermore, state-of-the-art monitoring systems for ship movements in port waters, along with non-intrusive cargo inspection methods like X-ray scanners and RFID tracking for cargo integrity, support prompt response times to security incidents [27]. The integrity, safety, and operational continuity of port infrastructures—which are crucial hubs in both national security frameworks and international trade networks—must be preserved by the integration of these security measures.
II-C Supply Chain Security
@@
Supply chain security refers to the laws, practices, and technological advancements created to protect physical items, information flows, and supply chain assets against threats such as theft, terrorism, and natural catastrophes[29]. The process of moving commodities by sea, involving shipping, port operations, and frequently interior transportation, is referred to as a maritime supply chain. This intricate system connects numerous ports spread over several nations and continents and encompasses a wide spectrum of stakeholders, including freight forwarders, shippers, carriers, and customs officials. The maritime supply chain is faced with a number of difficulties that could seriously jeopardize operational effectiveness and security. Among the many potential reasons of port interruptions, there include worker strikes, natural catastrophes, equipment malfunctions, and cyberattacks [30]. Port disruptions are considered a critical vulnerability. Information security threats, such as cyberattacks on logistics management systems and data breaches revealing private shipping information, pose a significant concern as well [31]. Physical security threats to marine supply chains include piracy (in high-risk maritime zones), cargo theft, and the smuggling of illegal commodities [29]. These are problems that have not yet been properly resolved. As a result of continued climate change and rising sea levels, port infrastructure and shipping routes are becoming more and more susceptible to catastrophic weather events [32]. Furthermore, operational risks such as equipment failures, delays in cargo handling, and human mistake in logistics planning can have a major impact on the efficiency of supply chains. Considering that over 80% of worldwide trade is carried out by ships [33], any interruption to the movement of products around the world would be catastrophic.
@@
The suppliers’ perspective is a crucial aspect of the marine supply chain, particularly with regard to the tools and equipment they provide that are essential to numerous nautical systems. This aspect of supply chain security is vital and warrants detailed examination to fully understand the scope of maritime supply chain security. The maritime sector includes a range of specialized technology, including cargo handling devices, propulsion systems, communication tools, and numerous sensors. A breach in one or more product supply chains might lead to many vulnerabilities in important maritime systems. Malicious actors attempting to install hardware or software intended to interfere with operations and obtain unauthorized access may take advantage of these vulnerabilities [34]. On the other hand, it may lead to the replacement of original parts with fake or inferior ones, which would raise the danger of equipment malfunctions and security issues [29]. Moreover, the introduction of malicious code into software or firmware during the manufacturing process may result in the creation of a backdoor that allows for unauthorized control [30].
II-D Risk Assessment Method